# Duality theory in linear optimization and its extensions — formally verified

### v2

@@ 2.1 Review of algebraic typeclasses that our project depends on
zero_add : $\forall$ a : M, 0 + a = a
add_zero : $\forall$ a : M, a + 0 = a
@@
nsmul : $\mathbb{N}$rightarrow M $\rightarrow$ M
nsmul_zero : $\forall$ x : M, nsmul 0 x = 0
nsmul_succ : $\forall$ (n : $\mathbb{N}$) (x : M), nsmul (n + 1) x = nsmul n x + x
@@
npow : $\mathbb{N}$rightarrow M $\rightarrow$ M
npow_zero : $\forall$ x : M, npow 0 x = 1
npow_succ : $\forall$ (n : $\mathbb{N}$) (x : M), npow (n + 1) x = npow n x * x
@@
sub_eq_add_neg : $\forall$ a b : G, a - b = a + -b
zsmul : $\mathbb{Z}$rightarrow G $\rightarrow$ G
zsmul_zero : $\forall$ a : G, zsmul 0 a = 0
zsmul_succ (n : $\mathbb{N}$) (a : G) : zsmul n.succ a = zsmul n a + a
zsmul_neg (n : $\mathbb{N}$) (a : G) : zsmul (Int.negSucc n) a = -(zsmul n.succ a)
@@
div_eq_mul_inv : $\forall$ a b : G, a / b = a * b-1
zpow : $\mathbb{Z}$rightarrow G $\rightarrow$ G
zpow_zero : $\forall$ a : G, zpow 0 a = 1
zpow_succ (n : $\mathbb{N}$) (a : G) : zpow n.succ a = zpow n a * a
zpow_neg (n : $\mathbb{N}$) (a : G) : zpow (Int.negSucc n) a = (zpow n.succ a)-1
@@
left_distrib : $\forall$ a b c : R, a * (b + c) = a * b + a * c
right_distrib : $\forall$ a b c : R, (a + b) * c = a * c + b * c
@@
class MulZeroClass (M0 : Type u) extends Mul M0, Zero M0 where
zero_mul : $\forall$ a : M0, 0 * a = 0
mul_zero : $\forall$ a : M0, a * 0 = 0
class NonUnitalNonAssocSemiring ($\alpha$ : Type u) extends AddCommMonoid $\alpha$, Distrib $\alpha$, MulZeroClass $\alpha$
@@
class SemigroupWithZero (S0 : Type u) extends Semigroup S0, MulZeroClass S0
class NonUnitalSemiring ($\alpha$ : Type u) extends NonUnitalNonAssocSemiring $\alpha$, SemigroupWithZero $\alpha$
@@
natCast_succ : $\forall$ n : $\mathbb{N}$, natCast (n + 1) = (natCast n) + 1
@@
intCast_ofNat : $\forall$ n : $\mathbb{N}$, intCast (n : $\mathbb{N}$) = Nat.cast n
intCast_negSucc : $\forall$ n : $\mathbb{N}$, intCast (Int.negSucc n) = - Nat.cast (n + 1)
@@
class MulZeroOneClass (M0 : Type u) extends MulOneClass M0, MulZeroClass M0
class NonAssocSemiring ($\alpha$ : Type u) extends NonUnitalNonAssocSemiring $\alpha$, MulZeroOneClass $\alpha$,
AddCommMonoidWithOne $\alpha$
@@
class MonoidWithZero (M0 : Type u) extends Monoid M0, MulZeroOneClass M0, SemigroupWithZero M0
class Semiring ($\alpha$ : Type u) extends NonUnitalSemiring $\alpha$, NonAssocSemiring $\alpha$, MonoidWithZero $\alpha$
@@
class CommRing ($\alpha$ : Type u) extends Ring $\alpha$, CommMonoid $\alpha$
Division ring is a nontrivial ring whose multiplication forms a division monoid, whose nonzero elements have multiplicative inverses, whose zero is inverse to itself (if you find the equality $00^{-1}\!=\!$ 0 disturbing, read [25] that explains it), and embeds rational numbers:
class Nontrivial ($\alpha$ : Type*) : Prop where
exists_pair_ne : $\exists$ x y : $\alpha$, x $\neq$ y
@@
mul_inv_cancel : $\forall$ (a : K), a $\neq$ 0 $\rightarrow$ a * a-1 = 1
inv_zero : (0 : K)-1 = 0
@@
Preorder is a reflexive & transitive relation on any structure with binary relational symbols $\leq$ and $<$ where the strict comparison $a<b$ is equivalent to $a\leq b\operatorname{\;\wedge\;}\neg(b\leq a)$ given by the relation $\leq$ which is neither required to be symmetric nor required to be antisymmetric:
class Preorder ($\alpha$ : Type u) extends LE $\alpha$, LT $\alpha$ where
le_refl : $\forall$ a : $\alpha$, a $\leq$ a
le_trans : $\forall$ a b c : $\alpha$, a $\leq$ b $\rightarrow$ b $\leq$ c $\rightarrow$ a $\leq$ c
lt_iff_le_not_le : $\forall$ a b : $\alpha$, a < b $\iff$ a $\leq$ b $\wedge$ $\neg$¬b $\leq$ a
@@
class PartialOrder ($\alpha$ : Type u) extends Preorder $\alpha$ where
le_antisymm : $\forall$ a b : $\alpha$, a $\leq$ b $\rightarrow$ b $\leq$ a $\rightarrow$ a = b
@@
class OrderedAddCommGroup ($\alpha$ : Type u) extends AddCommGroup $\alpha$, PartialOrder $\alpha$ where
add_le_add_left : $\forall$ a b : $\alpha$, a $\leq$ b $\rightarrow$ $\forall$ c : $\alpha$, c + a $\leq$ c + b
@@
class StrictOrderedRing ($\alpha$ : Type u) extends Ring $\alpha$, OrderedAddCommGroup $\alpha$, Nontrivial $\alpha$ where
zero_le_one : 0 $\leq$ (1 : $\alpha$)
mul_pos : $\forall$ a b : $\alpha$, 0 < a $\rightarrow$ 0 < b $\rightarrow$ 0 < a * b
@@
class LinearOrder ($\alpha$ : Type u) extends PartialOrder $\alpha$, (...)
le_total (a b : $\alpha$) : a $\leq$ b $\vee$ b $\leq$ a
@@
class LinearOrderedAddCommGroup ($\alpha$ : Type u) extends OrderedAddCommGroup $\alpha$, LinearOrder $\alpha$
@@
class LinearOrderedRing ($\alpha$ : Type u) extends StrictOrderedRing $\alpha$, LinearOrder $\alpha$
@@
class LinearOrderedCommRing ($\alpha$ : Type u) extends LinearOrderedRing $\alpha$, CommMonoid $\alpha$
@@
class LinearOrderedField ($\alpha$ : Type*) extends LinearOrderedCommRing $\alpha$, Field $\alpha$
Note that LinearOrderedDivisionRing is not a part of the algebraic hierarchy provided by Mathlib, hence LinearOrderedField does not inherit LinearOrderedDivisionRing. Because of that, we provide a custom instance that converts LinearOrderedField to LinearOrderedDivisionRing as follows:
@@ 2.2 Extended linearly ordered fields
@[coe] def toE : F $\rightarrow$ (Extend F) := some $\circ$ some
[…]
@@
  • negating any finite value $x$ gives $-x$ converted to the type $F_{\infty}$
@@
  • multiplying $\top$ by 00 from the left gives 00; multiplying $\top$ by a strictly positive element from the left gives $\top$
@@ 2.3 Vectors and matrices
$\begin{pmatrix}1&2&3\\ 4&5&6\\ \end{pmatrix}\cdot\begin{pmatrix}7\\ 8\\ 9\end{pmatrix}=\begin{pmatrix}?\\ \_\end{pmatrix}$
@@
One thing to be said about the implementation of matrices is that we lied a little bit when we wrote things like $A:(I\times J)\rightarrow F$. In reality, A : Matrix I J F is implemented as $A:\,I\rightarrow(J\rightarrow F)$. This technique [27] is called currying and makes it easier to pass individual rows of the matrix into functions and lemmas.
The following instance defines how explicit vectors are compared; Mathlib defines $x\leq y$ to hold if and only if $x_{i}\leq y_{i}$ holds for every $i$ (it is, in fact, more general than we said because the following instance talks about $\Pi$-types [19], i.e., the type of x i can depend on the value of i):
instance Pi.hasLe $\iota$ : Type*} $\pi$ : $\iota$rightarrow Type*} [$\forall$ i, LE ($\pi$ i)] :
LE ($\forall$ i, $\pi$ i) where le x y := $\forall$ i, x i $\leq$ y i
@@ 2.3.1 Mathlib definitions versus our definitions
def dotProduct [Fintype m] [Mul $\alpha$] [AddCommMonoid $\alpha$] (v w : m $\rightarrow$ $\alpha$) : $\alpha$ :=
$\sum$ i, v i * w i
@@
def Matrix.mulVec [Fintype n] [NonUnitalNonAssocSemiring $\alpha$] (M : Matrix m n $\alpha$) (v : n $\rightarrow$ $\alpha$) : m $\rightarrow$ $\alpha$
[…]
@@
[…]
scoped infixr:73 " *v " => Matrix.mulVec
@@
variable $\alpha$ $\gamma$ : Type*} [AddCommMonoid $\alpha$] [SMul $\gamma$ $\alpha$]
def dotWeig (v : I $\rightarrow$ $\alpha$) (w : I $\rightarrow$ $\gamma$) : $\alpha$ :=
$\sum$ i : I, w i $\bullet$ v i
infixl:72 " $\cdot$ " => dotWeig
def Matrix.mulWeig (M : Matrix I J $\alpha$) (w : J $\rightarrow$ $\gamma$) (i : I) : $\alpha$ :=
M i $\cdot$ w
infixr:73 " m* " => Matrix.mulWeig
We start by declaring that $\alpha$ and $\gamma$ are types from any universe (not necessarily both from the same universe). We require that $\alpha$ forms an abelian monoid and that $\gamma$ has a scalar action on $\alpha$. In this setting, we can instantiate $\alpha$ with $F_{\infty}$ and $\gamma$ with $F$ for any linearly ordered field $F$.
@@ 2.4 Modules and how to order them
class MulAction ($\alpha$ : Type*) ($\beta$ : Type*) [Monoid $\alpha$] extends SMul $\alpha$ $\beta$ where
one_smul : $\forall$ b : $\beta$, (1 : $\alpha$) $\bullet$ b = b
mul_smul : $\forall$ (x y : $\alpha$) (b : $\beta$), (x * y) $\bullet$ b = x $\bullet$ y $\bullet$ b
@@
smul_zero : $\forall$ a : M, a $\bullet$ (0 : A) = 0
smul_add : $\forall$ (a : M) (x y : A), a $\bullet$ (x + y) = a $\bullet$ x + a $\bullet$ y
@@
add_smul : $\forall$ (r s : R) (x : M), (r + s) $\bullet$ x = r $\bullet$ x + s $\bullet$ x
zero_smul : $\forall$ x : M, (0 : R) $\bullet$ x = 0
Note the class Module does not extend the class Semiring; instead, it requires Semiring as an argument. The abelian monoid is also required as argument in the definition. We call such a class “mixin”. Thanks to this design, we do not need to define superclasses of Module in order to require “more than a module”. Instead, we use superclasses in the respective arguments, i.e., “more than a semiring” and/or “more than an abelian monoid”. For example, if we replace
@@
[LinearOrderedDivisionRing R] [AddCommGroup W] [Module R W] [LinearOrderedAddCommGroup V] [Module R V]
@@
class PosSMulMono ($\alpha$ $\beta$ : Type*) [SMul $\alpha$ $\beta$] [Preorder $\alpha$] [Preorder $\beta$] [Zero $\alpha$] : Prop where
[…]
@@ 2.5 Linear programming
[…]
@@
hAi : $\neg$¬$\exists$ i : I, ($\exists$ j : J, A i j = $\bot$) $\wedge$ ($\exists$ j : J, A i j = $\top$)
hAj : $\neg$¬$\exists$ j : J, ($\exists$ i : I, A i j = $\bot$) $\wedge$ ($\exists$ i : I, A i j = $\top$)
hbA : $\neg$¬$\exists$ i : I, ($\exists$ j : J, A i j = $\bot$) $\wedge$ b i = $\bot$
hcA : $\neg$¬$\exists$ j : J, ($\exists$ i : I, A i j = $\top$) $\wedge$ c j = $\bot$
hAb : $\neg$¬$\exists$ i : I, ($\exists$ j : J, A i j = $\top$) $\wedge$ b i = $\top$
hAc : $\neg$¬$\exists$ j : J, ($\exists$ i : I, A i j = $\bot$) $\wedge$ c j = $\top$
The following definition says how linear programs are dualized (first, the abbreviation ExtendedLP.dualize says that $(A,b,c)$ is mapped to $(-A^{T},c,b)$; then the definition ValidELP.dualize says that the dualization of ValidELP is inherited from ExtendedLP.dualize; the remaining six lines generate proofs that our six conditions stay satisfied after dualization, where aeply is a custom tactic based on aesop [17] and apply):
@@
[…]
if $\neg$¬P.IsFeasible then
some $\top$
@@
if hr : $\exists$ r : F, P.Reaches (toE r) $\wedge$ P.IsBoundedBy r then
@@
For example, OppositesOpt (-3) 3 and OppositesOpt 5 (-5) hold. The last line says that OppositesOpt none none is false. We later show that none is never the optimum of a valid222 In principle, we could say that none is never the optimum of any linear program, i.e., $P^{\ast}$ is defined for every (extended) linear program $P$, but we did not bother proving it. linear program, but we do not know it yet.

### v3

@@ 2.1 Review of algebraic typeclasses that our project depends on
zero_add : a : M, 0 + a = a
add_zero : a : M, a + 0 = a
@@
nsmul : M M
nsmul_zero : x : M, nsmul 0 x = 0
nsmul_succ : (n : ℕ) (x : M), nsmul (n + 1) x = nsmul n x + x
@@
npow : M M
npow_zero : x : M, npow 0 x = 1
npow_succ : (n : ℕ) (x : M), npow (n + 1) x = npow n x * x
@@
sub_eq_add_neg : a b : G, a - b = a + -b
zsmul : G G
zsmul_zero : a : G, zsmul 0 a = 0
zsmul_succ (n : ℕ) (a : G) : zsmul n.succ a = zsmul n a + a
zsmul_neg (n : ℕ) (a : G) : zsmul (Int.negSucc n) a = -(zsmul n.succ a)
@@
div_eq_mul_inv : a b : G, a / b = a * b⁻¹
zpow : G G
zpow_zero : a : G, zpow 0 a = 1
zpow_succ (n : ℕ) (a : G) : zpow n.succ a = zpow n a * a
zpow_neg (n : ℕ) (a : G) : zpow (Int.negSucc n) a = (zpow n.succ a)⁻¹
@@
left_distrib : a b c : R, a * (b + c) = a * b + a * c
right_distrib : a b c : R, (a + b) * c = a * c + b * c
@@
class MulZeroClass (M : Type u) extends Mul M₀, Zero M where
zero_mul : a : M₀, 0 * a = 0
mul_zero : a : M₀, a * 0 = 0
class NonUnitalNonAssocSemiring : Type u) extends AddCommMonoid α, Distrib α, MulZeroClass α
@@
class SemigroupWithZero (S : Type u) extends Semigroup S₀, MulZeroClass S
class NonUnitalSemiring : Type u) extends NonUnitalNonAssocSemiring α, SemigroupWithZero α
@@
natCast_succ : n : ℕ, natCast (n + 1) = (natCast n) + 1
@@
intCast_ofNat : n : ℕ, intCast (n : ℕ) = Nat.cast n
intCast_negSucc : n : ℕ, intCast (Int.negSucc n) = - Nat.cast (n + 1)
@@
class MulZeroOneClass (M : Type u) extends MulOneClass M₀, MulZeroClass M
class NonAssocSemiring : Type u) extends NonUnitalNonAssocSemiring α, MulZeroOneClass α,
AddCommMonoidWithOne α
@@
class MonoidWithZero (M : Type u) extends Monoid M₀, MulZeroOneClass M₀, SemigroupWithZero M
class Semiring : Type u) extends NonUnitalSemiring α, NonAssocSemiring α, MonoidWithZero α
@@
class CommRing : Type u) extends Ring α, CommMonoid α
Division ring is a nontrivial ring whose multiplication forms a division monoid, whose nonzero elements have multiplicative inverses, whose zero is inverse to itself (if you find the equality $0^{-1}\!=\!0$ disturbing, read [DivBy0] that explains it), and embeds rational numbers:
class Nontrivial : Type*) : Prop where
exists_pair_ne : x y : α, x y
@@
mul_inv_cancel : (a : K), a 0 a * a⁻¹ = 1
inv_zero : (0 : K)⁻¹ = 0
@@
Preorder is a reflexive & transitive relation on any structure with binary relational symbols $\leq$ and << where the strict comparison $a<b$ is equivalent to $a\leq b\operatorname{\;\wedge\;}\neg(b\leq a)$ given by the relation $\leq$ which is neither required to be symmetric nor required to be antisymmetric:
class Preorder : Type u) extends LE α, LT α where
le_refl : a : α, a a
le_trans : a b c : α, a b b c a c
lt_iff_le_not_le : a b : α, a < b a b ¬b a
@@
class PartialOrder : Type u) extends Preorder α where
le_antisymm : a b : α, a b b a a = b
@@
class OrderedAddCommGroup : Type u) extends AddCommGroup α, PartialOrder α where
add_le_add_left : a b : α, a b c : α, c + a c + b
@@
class StrictOrderedRing : Type u) extends Ring α, OrderedAddCommGroup α, Nontrivial α where
zero_le_one : 0 (1 : α)
mul_pos : a b : α, 0 < a 0 < b 0 < a * b
@@
class LinearOrder : Type u) extends PartialOrder α, (…)
le_total (a b : α) : a b b a
@@
class LinearOrderedAddCommGroup : Type u) extends OrderedAddCommGroup α, LinearOrder α
@@
class LinearOrderedRing : Type u) extends StrictOrderedRing α, LinearOrder α
@@
class LinearOrderedCommRing : Type u) extends LinearOrderedRing α, CommMonoid α
@@
class LinearOrderedField : Type*) extends LinearOrderedCommRing α, Field α
Note that LinearOrderedDivisionRing is not a part of the algebraic hierarchy provided by Mathlib, hence LinearOrderedField does not inherit LinearOrderedDivisionRing.
Because of that, we provide a custom instance that converts LinearOrderedField to LinearOrderedDivisionRing as follows:
@@ 2.2 Extended linearly ordered fields
@[coe] def toE : F (Extend F) := some some
instance : Coe F (Extend F) := toE
@@
  • negating any finite value $x$ gives x-x converted to the type $F_{\infty}$
@@
  • multiplying $\top$ by 0 from the left gives 0; multiplying $\top$ by a strictly positive element from the left gives $\top$
@@ 2.3 Vectors and matrices
[…]
@@
One thing to be said about the implementation of matrices is that we lied a little bit when we wrote things like $A:(I\times J)\rightarrow F$. In reality, A : Matrix I J F is implemented as $A:\,I\rightarrow(J\rightarrow F)$. This technique [Currying] is called currying and makes it easier to pass individual rows of the matrix into functions and lemmas.
The following instance defines how explicit vectors are compared; Mathlib defines $x\leq y$ to hold if and only if $x_{i}\leq y_{i}$ holds for every $i$ (it is, in fact, more general than we said because the following instance talks about $\Pi$-types [MLTT], i.e., the type of x i can depend on the value of i):
instance Pi.hasLe : Type*} : ι Type*} [∀ i, LE i)] :
LE (∀ i, π i) where le x y := i, x i y i
@@ 2.3.1 Mathlib definitions versus our definitions
def dotProduct [Fintype m] [Mul α] [AddCommMonoid α] (v w : m α) : α :=
i, v i * w i
@@
def Matrix.mulVec [Fintype n] [NonUnitalNonAssocSemiring α] (M : Matrix m n α) (v : n α) : m α
| i => (fun j => M i j) ⬝ᵥ v
@@
infixl:72 " ⬝ᵥ " => dotProduct
scoped infixr:73 " *ᵥ " => Matrix.mulVec
@@
variable γ : Type*} [AddCommMonoid α] [SMul γ α]
def dotWeig (v : I α) (w : I γ) : α :=
i : I, w i v i
infixl:72 " ᵥ⬝ " => dotWeig
def Matrix.mulWeig (M : Matrix I J α) (w : J γ) (i : I) : α :=
M i ᵥ⬝ w
infixr:73 " ₘ* " => Matrix.mulWeig
We start by declaring that $\alpha$ and $\gamma$ are types from any universe (not necessarily both from the same universe). We require that $\alpha$ forms an abelian monoid and that $\gamma$ has a scalar action on $\alpha$. In this setting, we can instantiate $\alpha$ with $F_{\infty}$ and $\gamma$ with $F_{\geq 0}$ for any linearly ordered field $F$.
@@ 2.4 Modules and how to order them
class MulAction : Type*) : Type*) [Monoid α] extends SMul α β where
one_smul : b : β, (1 : α) b = b
mul_smul : (x y : α) (b : β), (x * y) b = x y b
@@
smul_zero : a : M, a (0 : A) = 0
smul_add : (a : M) (x y : A), a (x + y) = a x + a y
@@
add_smul : (r s : R) (x : M), (r + s) x = r x + s x
zero_smul : x : M, (0 : R) x = 0
Note the class Module does not extend the class Semiring; instead, it requires Semiring as an argument. The abelian monoid is also required as argument in the definition. We call such a class “mixin”. Thanks to this design, we do not need to define subclasses of Module in order to require “more than a module”. Instead, we use subclasses in the respective arguments, i.e., “more than a semiring” and/or “more than an abelian monoid”. For example, if we replace
@@
[LinearOrderedDivisionRing R] [AddCommGroup W] [Module R W]
[LinearOrderedAddCommGroup V] [Module R V]
@@
class PosSMulMono β : Type*) [SMul α β] [Preorder α] [Preorder β] [Zero α] : Prop where
elim {a : α} (ha : 0 a) {b b : β} (hb : b b₂) : a b a b
@@ 2.5 Linear programming
def ExtendedLP.Reaches [Fintype J] (P : ExtendedLP I J F) (r : F∞) : Prop :=
x : J F≥0, P.IsSolution x P.c ᵥ⬝ x = r
@@
hAi : ¬∃ i : I, (∃ j : J, A i j = ⊥) (∃ j : J, A i j = ⊤)
hAj : ¬∃ j : J, (∃ i : I, A i j = ⊥) (∃ i : I, A i j = ⊤)
hbA : ¬∃ i : I, (∃ j : J, A i j = ⊥) b i =
hcA : ¬∃ j : J, (∃ i : I, A i j = ⊤) c j =
hAb : ¬∃ i : I, (∃ j : J, A i j = ⊤) b i =
hAc : ¬∃ j : J, (∃ i : I, A i j = ⊥) c j =
The following definition says how linear programs are dualized (first, the abbreviation ExtendedLP.dualize says that $(A,b,c)$ is mapped to $(-A^{T},c,b)$; then ValidELP.dualize says that the dualization of ValidELP is inherited from ExtendedLP.dualize; the remaining six lines generate proofs that our six conditions stay satisfied after dualization, where aeply is a custom tactic based on aesop [Aesop] and apply):
@@
noncomputable def ExtendedLP.optimum [Fintype J] (P : ExtendedLP I J F) : Option F :=
if ¬P.IsFeasible then
some
@@
if hr : r : F, P.Reaches (toE r) P.IsBoundedBy r then
@@
For example, OppositesOpt (-3) 3 and OppositesOpt 5 (-5) hold. The last line says that OppositesOpt none none is false. We later show that none is never the optimum of a valid [^4] In principle, we could say that none is never the optimum of any linear program, i.e., $P^{\ast}$ is defined for every (extended) linear program $P$, but we did not bother proving it. linear program, but we do not know it yet.